I need help with this question.
f(x)=2x+3 and g(x)=4-x^(2);
a) (f*g)(5)
b) (f*f)(5)
c) (g*g)(-1)
d) (g*f)(-5)
I got this completely wrong and I would like to see where I may have gotten it mixed up.
Here are the answers I got...
a) -273
b) 169
c) 9
d) 147

Answer by greenestamps(13367)



Your answers are fine.... Why do you think you got them all wrong?

(f*g)(5)=f(5)*g(5)=13(-21)=-273

etc....

--------------------------------------------------------------------

If your answers are marked wrong, then you need to ask your instructor why....